# Why do Machine Learning Notebooks Crash? An Empirical Study on Public Python Jupyter Notebooks

### v2

@@ 1 Introduction
With the growing availability of data, machine learning (ML) and deep learning (DL) techniques have gained popularity for data modeling and decision-making within the broader field of data science [4, 5]. In this context, Jupyter notebooks, particularly Python-based ones, have been widely adopted for ML prototyping and data analysis [6]. In industrial settings [7], Jupyter notebooks are often the starting point for developing ML applications before transitioning to production-ready ML pipelines, where much of the Python code used in notebooks is reused in large-scale distributed computing environments. This widespread adoption, even in mission-critical applications such as those at NASA Jet Propulsion Lab [8], underscores the importance of quality assurance for notebooks. Bugs that arise during the notebook stage may persist into later development stages, potentially affecting deployed products. Therefore, studying notebook bugs is essential to improve the overall reliability and robustness of ML-driven software.
@@ 4.1 RQ1: What are the predominant crash types and root causes of crashes in ML notebooks?
To assess the accuracy of our root cause analysis, we try to reproduce 30 crashes in Kaggle and successfully reproduce 20. We then resolve the crashes, finding that 18 ($90\%$) are caused by the root causes identified in this study. This high accuracy suggests that our approach is reliable.
@@ 6.1 External validity
As our dataset only includes publicly shared GitHub and Kaggle notebooks, it may exclude crashes that users fix before publishing, thus crashes encountered during development may be underrepresented. As a result, our findings reflect crashes that persist, which may not be the full spectrum of debugging challenges. Future work could address this by analyzing version histories or autosaved checkpoints to capture a more complete development process.
@@ 6.2 Internal validity
We manually label ML notebook crashes, which can be prone to errors. To enhance the classification reliability, we use a structured two-stage process. In the labeling stage, three evaluators independently and iteratively categorize crashes, followed by discussions to synchronize findings, reduce bias, and promote consistency. In the reviewing stage, a reviewer validates the labels. Any disagreements are resolved through consensus.
@@
The error outputs we analyze represent real execution crashes, but they are not necessarily reproducible due to factors like execution order, missing datasets, or environment replication [2]. While reproducing crashes could further validate our findings, it is challenging, especially for GitHub notebooks. Kaggle notebooks, with their accessible datasets and pre-configured environments, may enable reproducibility. Nevertheless, we carried out a small-scale validation to assess the accuracy of our root cause analysis by trying to reproduce a few crashes in Kaggle, and the high accuracy suggests that our approach is reliable (see Section 4.1). In future work, we plan to extend crash reproducibility efforts in Kaggle notebooks to validate our results thoroughly and create a reproducible dataset for deeper analysis.

### v3

@@ 1 Introduction
With the growing availability of data, machine learning (ML) and deep learning (DL) techniques have gained popularity for data modeling and decision-making within the broader field of data science [4, 5]. In this context, Jupyter notebooks, particularly Python-based ones, have been widely adopted for ML prototyping and data analysis [6]. As confirmed by several industrial collaborators [7], Jupyter notebooks often serve as the starting point for developing ML applications before transitioning to production-ready ML pipelines, where much of the Python code used in notebooks is reused in large-scale distributed computing environments. This widespread adoption, even in mission-critical applications such as those at NASA Jet Propulsion Lab [8], underscores the importance of quality assurance for notebooks. Bugs that arise during the notebook stage may persist into later development stages, potentially affecting deployed products. Therefore, studying notebook bugs is essential to improve the overall reliability and robustness of ML-driven software.
@@ 4.1 RQ1: What are the predominant crash types and root causes of crashes in ML notebooks?
Evaluating root cause accuracy
To assess the accuracy of our root cause analysis, we randomly selected 30 crashes in Kaggle notebooks and attempted to reproduce and fix them. For each case, we download the input datasets (if available), execute a subset of code cells necessary to trigger the crash, localize and repair the faulty code to verify if the labeled root cause aligns with the actual fault. A crash is considered fixed if the previously crashing cell executes successfully while preserving the original intent.
We successfully reproduced and fixed 20 out of the 30 crashes. Of the 10 unsuccessful cases, 8 lack input datasets, and the remaining two no longer crash. From the 20 cases, 18 ($90\%$) match the root cause identified by this study, which suggests that our approach is reliable. For the two misclassifications, one involves an index error due to out-of-order execution mislabeled as API misuse. The other is a tensor shape mismatch caused by faulty logic (i.e., implementation error) in earlier cells, incorrectly labeled as data confusion. The main challenge in labeling errors and root causes without executing the notebooks lies in mentally tracking the internal kernel state. This can lead to misunderstandings by labelers, especially when the code logic is complex.
@@ 6.1 External validity
As our dataset only includes publicly shared GitHub and Kaggle notebooks, it may exclude crashes that users fix before publishing, thus crashes encountered during development may be underrepresented. As a result, our findings reflect crashes that persist, which may not be the full spectrum of debugging challenges. Future work could address this by analyzing version histories or autosaved checkpoints to capture a more complete development process. Furthermore, although we have filtered out tutorial notebooks, the ML notebooks in our dataset may still not fully represent those used as prototypes in an industrial setting. This introduces a potential threat to validity when generalizing crash patterns that could propagate from notebooks to production. Future work should address this by studying ML prototype notebooks developed within industry to better assess how our findings apply.
@@ 6.2 Internal validity
We manually label ML notebook crashes, which can be prone to errors. To enhance the reliability of classification, we use a structured two-stage process. In the labeling stage, three evaluators independently and iteratively categorize crashes, followed by discussions to synchronize findings, reduce bias, and promote consistency. In the reviewing stage, a reviewer validates the labels. Any disagreements are resolved by consensus.
@@
The error outputs we analyze represent real execution crashes, but they are not necessarily reproducible due to factors such as execution order, missing datasets, or environment replication [2]. While reproducing crashes could further validate our findings, it is challenging, especially for GitHub notebooks. Kaggle notebooks, with their accessible datasets and pre-configured environments, may enable reproducibility. Nevertheless, we carried out a small-scale validation to assess the accuracy of our root cause analysis by trying to reproduce a few crashes in Kaggle, and the high accuracy suggests that our approach is reliable (see Section 4.1). In future work, we plan to extend crash reproducibility efforts in Kaggle notebooks to validate our results thoroughly and create a reproducible dataset for deeper analysis.